# These $\altaffiliationThese authors contributed equally to this work.$

### v1

@@ 13 Conclusion
We introduced the augmented MOLOPT family of all-electron Gaussian basis sets optimized for accurate excited-state calculations of large molecules and solids for the elements H to Cl. These basis sets achieve fast basis set convergence of $GW$ quasiparticle energy differences and BSE excitation energies while ensuring low condition numbers of the overlap matrix $\mathbf{S}$, thereby enabling numerically stable calculations. For $G_{0}W_{0}$@PBE0 gaps, aug-DZVP-MOLOPT-ae yields a mean absolute deviation (MAD) of 60 meV compared to the aug-cc-pV5Z complete basis set, outperforming the larger aug-cc-pVTZ basis set (MAD: 80 meV) for organic molecules. Similar MAD are observed for BSE and TDDFT excitation energies. The augmented MOLOPT basis sets exhibit excellent numerical stability, with overlap matrix condition numbers below $10^{7}$ even for 2000-atom nanographenes. We also generate very compact basis sets, aug-SZV-MOLOPT-ae-mini, which enable very efficient large-scale $G_{0}W_{0}$ calculations, e.g., on a 2312-atom nanographene ($L=128$) in just 3500 core hours and using only 2.9 TB of RAM. This demonstrates that the proposed augmented MOLOPT basis sets enable routine $GW$ and BSE calculations on large-scale systems with over 1000 atoms, keeping good numerical accuracy and reducing the computational cost by two orders of magnitude compared to previously used aug-cc-pVXZ basis sets. All generated augmented MOLOPT basis sets are freely available in the Supporting Information.
@@
Inputs and outputs of all calculations reported in this work are available in a Github repository 90.The augmented MOLOPT basis sets and corresponding auxiliary RI basis sets generated in this work are available in the supporting information S3, S4. The $GW$, GWGW+BSE and TDDFT algorithms employed in this work are available in the open-source package CP2K 57, 41.
@@
We thank Tilo Wettig for an important remark on recomputing sparse tensor elements, which led to the memory-saving scheme described in the appendix. We further thank Ritaj Tyagi for helpful comments on the basis sets and Momme Allalen, Gerald Mathias, Ferdinand Evers, Dorothea Golze, Jürg Hutter, Patrick Rinke, Carlo Pignedoli and Robert Schade for helpful discussions. We acknowledge funding by the Free State of Bavaria through the KONWIHR software initiative. DFG is acknowledged for funding via the Emmy Noether Programme (project number 503985532), CRC 1277 (project number 314695032, subproject A03) and RTG 2905 (project number 502572516). The authors gratefully acknowledge the computing time provided to them on the high performance computer Noctua 2 at the NHR Center PC2. These are funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for the national high-performance computing at universities (www.nhr-verein.de/unsere-partner). The authors also gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (www.lrz.de).
Appendix: Memory-saving scheme for low-scaling GW calculations
@@
In this appendix, we describe a memory saving scheme to reduce the random access memory (RAM) of low-scaling GW calculations 75, 27 substantially. The RAM bottleneck of the GW algorithm 75, 27 appears in the computation of the self-energy $\Sigma$ in imaginary time $i\tau$,
@@
The repeated calculation of 3cIs comes with the drawback that we need to compute the same integral $(\mu\nu|P)$ several times. Here, we make use of the properties of the Gaussian basis that analytical integral expressions are available for $(\mu\nu|P)$, such that this additional computational load is small 91. In fact, the computation of 3cIs for large systems exceeding 100 atoms only takes $0.1\,\%$ of the total execution time in the present GW algorithm 27.
@@
Table 3: Basis sizes, memory requirements for storing three-center integrals $(\mu\nu|P)$, Eq. (41), their sparsity, available RAM and the execution time for GWGW calculations on nanographenes from Fig. 6. The available memory for the storage of 3cIs is 6 GB per MPI rank throughout all the listed calculations. Compute nodes with larger memory capacity are used for system sizes $L\geq 32$ to accommodate the RAM needed for the DFT module to store memory-heavy plane-wave grids for computing the Hartree potential from Ewald summation. Execution time has been measured on the Noctua2 cluster (AMD Milan 7763) at PC2 computing center in Paderborn.
@@
1 14 12 aug-SZV-MOLOPT-ae-mini 198 2 11 178 100.00 % 0 1 240 0.001
1 14 12 aug-SZV-MOLOPT-ae-SR 268 5 18 312 100.00 % 0 1 240 0.003
1 14 12 aug-SZV-MOLOPT-ae 268 6 23 394 100.00 % 0 1 240 0.006
1 14 12 aug-DZVP-MOLOPT-ae 528 9 35 598 100.00 % 1 1 240 0.019
1 14 12 aug-$c$-pVDZ 430 23 72 1284 100.00 % 1 1 240 0.043
2 28 16 aug-SZV-MOLOPT-ae-mini 348 2 11 340 100.00 % 0 1 240 0.004
2 28 16 aug-SZV-MOLOPT-ae-SR 488 5 18 584 100.00 % 1 1 240 0.014
2 28 16 aug-SZV-MOLOPT-ae 488 6 23 740 100.00 % 1 1 240 0.027
2 28 16 aug-DZVP-MOLOPT-ae 984 9 35 1124 100.00 % 8 1 240 0.289
2 28 16 aug-$c$-pVDZ 788 23 72 2384 100.00 % 11 1 240 0.373
4 56 24 aug-SZV-MOLOPT-ae-mini 648 2 11 664 70.31 % 1 1 240 0.030
4 56 24 aug-SZV-MOLOPT-ae-SR 928 5 18 1128 77.07 % 5 1 240 0.141
4 56 24 aug-SZV-MOLOPT-ae 928 6 23 1432 91.05 % 8 1 240 0.366
4 56 24 aug-DZVP-MOLOPT-ae 1896 9 35 2176 99.82 % 62 1 240 4.204
4 56 24 aug-$c$-pVDZ 1504 23 72 4584 99.97 % 82 1 240 4.907
8 112 40 aug-SZV-MOLOPT-ae-mini 1248 2 11 1312 25.33 % 4 1 240 0.151
8 112 40 aug-SZV-MOLOPT-ae-SR 1808 5 18 2216 28.62 % 16 1 240 0.811
8 112 40 aug-SZV-MOLOPT-ae 1808 6 23 2816 38.33 % 28 1 240 2.143
8 112 40 aug-DZVP-MOLOPT-ae 3720 9 35 4280 50.68 % 240 1 240 33.487
8 112 40 aug-$c$-pVDZ 2936 23 72 8984 55.16 % 341 1 240 44.360
16 224 72 aug-SZV-MOLOPT-ae-mini 2448 2 11 2608 7.43 % 9 1 240 0.662
16 224 72 aug-SZV-MOLOPT-ae-SR 3568 5 18 4392 8.45 % 37 1 240 3.843
16 224 72 aug-SZV-MOLOPT-ae 3568 6 23 5584 11.73 % 66 1 240 9.099
16 224 72 aug-DZVP-MOLOPT-ae 7368 9 35 8488 16.15 % 595 10 2400 13.034
16 224 72 aug-$c$-pVDZ 5800 23 72 17784 18.01 % 861 10 2400 18.726
32 448 136 aug-SZV-MOLOPT-ae-mini 4848 2 11 5200 2.00 % 19 3 2850 0.808
32 448 136 aug-SZV-MOLOPT-ae-SR 7088 5 18 8744 2.28 % 80 1 240 13.650
32 448 136 aug-SZV-MOLOPT-ae 7088 6 23 11120 3.21 % 143 5 4750 6.116
64 896 264 aug-SZV-MOLOPT-ae-mini 9648 2 11 10384 0.52 % 40 3 2850 2.611
64 896 264 aug-SZV-MOLOPT-ae-SR 14128 5 18 17448 0.59 % 164 4 960 12.398
64 896 264 aug-SZV-MOLOPT-ae 14128 6 23 22192 0.84 % 297 5 4750 18.621
128 1792 520 aug-SZV-MOLOPT-ae-mini 19248 2 11 20752 0.13 % 81 3 2850 9.080
128 1792 520 aug-SZV-MOLOPT-ae-SR 28208 5 18 34856 0.15 % 334 4 3800 40.488
128 1792 520 aug-SZV-MOLOPT-ae 28208 6 23 44336 0.21 % 605 5 4750 65.225

### v2

@@ 5 Conclusion
We introduced the augmented MOLOPT family of all-electron Gaussian basis sets optimized for accurate excited-state calculations of large molecules for the elements H to Cl. These basis sets achieve fast basis set convergence of $GW$ quasiparticle energy differences and BSE excitation energies while ensuring low condition numbers of the overlap matrix $\mathbf{S}$, thereby enabling numerically stable calculations. For $G_{0}W_{0}$@PBE0 gaps, aug-DZVP-MOLOPT-ae yields a mean absolute deviation (MAD) of 60 meV compared to the aug-cc-pV5Z complete basis set, outperforming the larger aug-cc-pVTZ basis set (MAD: 80 meV) for organic molecules. Similar MAD are observed for BSE and TDDFT excitation energies. The augmented MOLOPT basis sets exhibit excellent numerical stability, with overlap matrix condition numbers below $10^{7}$ even for 9000-atom nanographenes. We also generate very compact basis sets, aug-SZV-MOLOPT-ae-mini, which enable very efficient large-scale $G_{0}W_{0}$ calculations , e.g., on a 9224-atom nanographene consuming only 34300 core hours. This demonstrates that the proposed augmented MOLOPT basis sets enable routine $GW$ and BSE calculations on large-scale systems with several thousands of atoms, keeping good numerical accuracy and reducing the computational cost by two orders of magnitude compared to previously used aug-cc-pVXZ basis sets. All generated augmented MOLOPT basis sets are freely available in the Supporting Information.
@@
Inputs and outputs of all calculations reported in this work are available in a Github repository github_repo and in a Zenodo database zenodo_repo.The augmented MOLOPT basis sets and corresponding auxiliary RI basis sets generated in this work are available in the supporting information S3, S4. The $GW$, GWGW+BSE and TDDFT algorithms employed in this work are available in the open-source package CP2K cp2k, Kuehne2020.
@@
We thank Tilo Wettig for an important remark on recomputing sparse tensor elements, which led to the memory-saving scheme described in the App. B. The support of Robert Schade and Xin Wu from the NHR Center Paderborn Center for Parallel Computing (PC2) with optimizing the $GW$ module in CP2K is gratefully acknowledged.
@@
We further thank Ritaj Tyagi for helpful comments on the basis sets and Momme Allalen, Gerald Mathias, Ferdinand Evers, Dorothea Golze, Jürg Hutter, Patrick Rinke and Carlo Pignedoli for helpful discussions. We acknowledge funding by the Free State of Bavaria through the KONWIHR software initiative. DFG is acknowledged for funding via the Emmy Noether Programme (project number 503985532), CRC 1277 (project number 314695032, subproject A03) and RTG 2905 (project number 502572516). The authors gratefully acknowledge the computing time made available to them on the high-performance computers Noctua2 Bauer2024 and Otus at the NHR Center Paderborn Center for Parallel Computing. This center is jointly supported by the Federal Ministry of Research, Technology and Space and the state governments participating in the National High-Performance Computing (NHR) joint funding program (www.nhr-verein.de/en/our-partners). The authors also gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (www.lrz.de).
Appendix App. A Computational details
Description of molecular test set
For benchmarking excited-state energies with our generated basis sets, we use the GWGW5000 dataset stuke2020atomic. We exclude all molecules with less than ten atoms as small molecules tend to have very diffuse unoccupied states; and the purpose of the generated basis sets is to describe large molecules with less diffuse unoccupied states. To reduce the computational cost, we only use molecules with at most 20 atoms. We also remove all molecules larger than 15 atoms in which carbon atoms outnumber all other non-hydrogen elements by more than a factor of two. Our aim is to ensure a balanced benchmark set avoiding overrepresentation of unsubstituted or weakly substituted hydrocarbons. The precise criterion for removal is $N_{\text{C}}\hskip 1.69998pt{>}\hskip 1.69998pt2(N_{\text{tot}}{-}N_{\text{H}}{-}N_{\text{C}}),$ where $N_{\text{C}}$ is the number of carbon atoms, $N_{\text{H}}$ the number of hydrogen atoms and $N_{\text{tot}}$ the total number of atoms in the molecule.
Applying these criteria gives 247 molecules in the GWGW5000 benchmark set, where the majority of the molecules contain C (98 %), H (96 %), N (76 %), O (74 %), while other elements are less often present: S (31 %), Cl (23 %), F (10 %), P (2 %), B (1 %) and Si (1 %). We also use a second molecular benchmark set that focuses on other elements (Li, Be, B, Na, Ca, Al, Si, P); we show the composition of this benchmark set and the calculations in the supplementary information SI1 and SI2, respectively. We employ the CP2K package for all calculations Kuehne2020, Kuehne2025. CP2K employs a Gaussian basis set for representing KS orbitals [Eq. (2)]. We use the Gaussian and augmented plane-waves scheme Lippert1999, which enables all-electron calculations in CP2K. We use implementations in CP2K of conventional $GW$ (Sec. 2.5) in imaginary-frequency formulation with analytic continuation wilhelm2016, BSE Graml2025 and TDDFT Hehn2022 (Sec. 2.6), as well as low-scaling $GW$ Wilhelm2021 (Sec. 3.24) based on the space-time method rojas_1995 using minimax time-frequency grids takatsuka2008minimax, hackbusch2019computation, Azizi2023, Azizi2024. We visualized atomic geometries using the VESTA program Momma2011.
Numerical aspects
We employ the PBE0 exchange-correlation functional Adamo1999 as starting point for our excited-state calculations. The usage of PBE0 as starting point for $GW$ and Bethe-Salpeter avoids numerical instabilities due to multipole features of the self-energy close to the quasiparticle solution Veril2018, Schambeck2024, which can be present when starting from the PBE functional Perdew1996. For the low–scaling $GW$ calculations on nanographenes (Sec. 4), however, we use PBE for the SCF cycle to reduce computational cost. As discussed in Ref.Schambeck2024, $G_{0}W_{0}$@PBE can suffer from numerical instabilities caused by poles in the self-energy $\Sigma_{n}(\omega)$ close to the quasiparticle energy $\omega\hskip 1.69998pt{\approx}\hskip 1.69998pt\varepsilon_{n}^{G_{0}W_{0}}$, where
$\varepsilon_{n}^{G_{0}W_{0}}=\varepsilon_{n}^{\text{PBE}}+\text{Re}\Sigma_{n}(\varepsilon_{n}^{G_{0}W_{0}})-v_{n}^{\text{xc}}.$ (32)
[…]
leading to the modified quasiparticle equation
Appendix App. B Memory-saving scheme for low-scaling GW calculations
@@
In this appendix, we describe a memory saving scheme to reduce the random access memory (RAM) of low-scaling $GW$ calculations Wilhelm2021, Graml2024 substantially. The RAM bottleneck of the GW algorithm Wilhelm2021, Graml2024 appears in the computation of the self-energy $\Sigma$ in imaginary time $i\tau$,
@@
The repeated calculation of 3cIs comes with the drawback that we need to compute the same integral $(\mu\nu|P)$ several times. Here, we make use of the properties of the Gaussian basis that analytical integral expressions are available for $(\mu\nu|P)$, such that this additional computational load is small libint. In fact, the computation of 3cIs for large systems exceeding 100 atoms only takes $0.1\,\%$ of the total execution time in the present GW algorithm Graml2024.
@@
Table 3: Basis sizes, memory requirements for storing three-center integrals $(\mu\nu|P)$, Eq. (40), their sparsity, available RAM and the execution time for GWGW calculations on nanographenes from Fig. 6. The available memory for the storage of 3cIs is 6 GB per MPI rank throughout all the listed calculations. Execution time has been measured on the Noctua2 cluster at PC2 computing center in Paderborn, where each node is equipped with two AMD Milan 7763 processors, each providing 64 cores (128 cores per node). For $L\geq 256$, the computations have been executed on the Otus cluster at the PC2 computing center, where one node consists of two AMD Turin 9655 processors, each providing 96 cores (192 cores per node).
@@
1 14 12 aug-SZV-MOLOPT-ae-mini 198 2 11 178 100.00 % 0 1 258 0.001
1 14 12 aug-SZV-MOLOPT-ae-SR 268 5 18 312 100.00 % 0 1 258 0.003
1 14 12 aug-SZV-MOLOPT-ae 268 6 23 394 100.00 % 0 1 258 0.006
1 14 12 aug-DZVP-MOLOPT-ae 528 9 35 598 100.00 % 1 1 258 0.019
1 14 12 aug-$c$-pVDZ 430 23 72 1284 100.00 % 1 1 258 0.043
2 28 16 aug-SZV-MOLOPT-ae-mini 348 2 11 340 100.00 % 0 1 258 0.004
2 28 16 aug-SZV-MOLOPT-ae-SR 488 5 18 584 100.00 % 1 1 258 0.014
2 28 16 aug-SZV-MOLOPT-ae 488 6 23 740 100.00 % 1 1 258 0.027
2 28 16 aug-DZVP-MOLOPT-ae 984 9 35 1124 100.00 % 8 1 258 0.289
2 28 16 aug-$c$-pVDZ 788 23 72 2384 100.00 % 11 1 258 0.373
4 56 24 aug-SZV-MOLOPT-ae-mini 648 2 11 664 70.31 % 1 1 258 0.030
4 56 24 aug-SZV-MOLOPT-ae-SR 928 5 18 1128 77.07 % 5 1 258 0.141
4 56 24 aug-SZV-MOLOPT-ae 928 6 23 1432 91.05 % 8 1 258 0.366
4 56 24 aug-DZVP-MOLOPT-ae 1896 9 35 2176 99.82 % 62 1 258 4.204
4 56 24 aug-$c$-pVDZ 1504 23 72 4584 99.97 % 82 1 258 4.907
8 112 40 aug-SZV-MOLOPT-ae-mini 1248 2 11 1312 25.33 % 4 1 258 0.151
8 112 40 aug-SZV-MOLOPT-ae-SR 1808 5 18 2216 28.62 % 16 1 258 0.811
8 112 40 aug-SZV-MOLOPT-ae 1808 6 23 2816 38.33 % 28 1 258 2.143
8 112 40 aug-DZVP-MOLOPT-ae 3720 9 35 4280 50.68 % 240 1 258 33.487
8 112 40 aug-$c$-pVDZ 2936 23 72 8984 55.16 % 341 1 258 44.360
16 224 72 aug-SZV-MOLOPT-ae-mini 2448 2 11 2608 7.43 % 9 1 258 0.662
16 224 72 aug-SZV-MOLOPT-ae-SR 3568 5 18 4392 8.45 % 37 1 258 3.843
16 224 72 aug-SZV-MOLOPT-ae 3568 6 23 5584 11.73 % 66 1 258 9.099
16 224 72 aug-DZVP-MOLOPT-ae 7368 9 35 8488 16.15 % 595 10 2577 13.034
16 224 72 aug-$c$-pVDZ 5800 23 72 17784 18.01 % 861 10 2577 18.726
32 448 136 aug-SZV-MOLOPT-ae-mini 4848 2 11 5200 2.00 % 19 3 3060 0.808
32 448 136 aug-SZV-MOLOPT-ae-SR 7088 5 18 8744 2.28 % 80 1 258 13.650
32 448 136 aug-SZV-MOLOPT-ae 7088 6 23 11120 3.21 % 143 5 5100 6.116
64 896 264 aug-SZV-MOLOPT-ae-mini 9648 2 11 10384 0.52 % 40 3 3060 2.611
64 896 264 aug-SZV-MOLOPT-ae-SR 14128 5 18 17448 0.59 % 164 4 1031 12.398
64 896 264 aug-SZV-MOLOPT-ae 14128 6 23 22192 0.84 % 297 5 5100 18.621
128 1792 520 aug-SZV-MOLOPT-ae-mini 19248 2 11 20752 0.13 % 81 3 3060 9.080
128 1792 520 aug-SZV-MOLOPT-ae-SR 28208 5 18 34856 0.15 % 334 4 4080 40.488
128 1792 520 aug-SZV-MOLOPT-ae 28208 6 23 44336 0.21 % 605 5 5100 65.225
256 3584 1032 aug-SZV-MOLOPT-ae-mini 38448 2 11 41488 0.03 % 163 10 16287 4.636
512 7168 2056 aug-SZV-MOLOPT-ae-mini 76848 2 11 82960 0.01 % 328 20 32573 8.928
Appendix App. C Basis set convergence  for 9,10-Dihydroanthracene
In this appendix, we report DFT and $GW$ HOMO-LUMO gaps as well as GWGW+BSE and TDDFT excitation energies of 9,10-Dihydroanthracene ($L\hskip 1.69998pt{=}\hskip 1.69998pt1$ nanographene in Sec. 4) using various basis sets to assess the quality of the basis sets as function of their basis set size. We plot in Fig. 9 the error with respect to the aug-cc-pV5Z calculation of the excitation energies and band gaps for the 9,10-Dihydroanthracene molecule as a function of the number of basis functions.
The results show similar trends as observed in Fig. 3. Both nonaugmented cc and MOLOPT basis sets show poor convergence with respect to the number of basis function, showing that these are not appropriate for the simulation of excited state energies. For the cc-pVQZ basis set, the error across all tests is around 150 meV on average, whereas it is around 450 meV on average for the comparable QZVPP-MOLOPT basis set, so that the cc basis sets perform better in this case (but still very poorly in comparison to the reference calculation, given the large size of the cc-pVQZ basis set).
The augmented basis sets show much better convergence across all four tests in Fig. 9 w.r.t. the basis set size. For the aug-cc-pVTZ basis set, which is of comparable size as the nonaugmented cc-pVQZ and QZVPP-MOLOPT basis sets, the error is around $20$ meV on average, and for the aug-TZVP-MOLOPT basis set it is around $10$ meV. Note that for the PBE0 HOMO-LUMO gap (Fig. 9a) and TDDFT excitation energies (Fig. 9d), the aug-cc-pVXZ basis sets converge faster with the basis set size than the aug-MOLOPT basis sets; still aug-MOLOPT basis sets feature improved numerical stability for large-scale calculations due to the reduced condition number compared to aug-cc-pVXZ. For the $GW$ HOMO-LUMO gap (Fig. 9b) and $GW$-BSE excitation energies (Fig. 9c), our aug-MOLOPT basis sets give faster convergence with the basis set size than aug-cc-pVXZ basis sets.
[…]